初級技術英語 1 単位 (選択)

Basic Technical English  カーペンターウォルター · 特任講師 / 機械工学科生産システム講座

【授業目的】 This course aims to increase the student’s ability to understand and

discuss professional materials in English. 【授業概要】 The general objectives of the course include: 1. Developing students’

reading and listening skills for academic, professional purposes 2. Assisting students’ ability to speak English in real-life situations 3. Helping students develop ability to understand English in various contexts. 4. Increase students’ knowledge of English vocabulary, especially words of a technical and scientific nature.

【キーワード】 Technical English， Vocabulary, Scientific 【先行科目】 [先行科目] 【関連科目】『中級技術英語』 (1.0) 【履修要件】 [要件] 【履修上の注意】 This course is a compulsory one for Nichia STC students. If you

are not a Nichia STC student and want to attend this class, please contact a coordinator in your department for getting permission for your attendance.

【到達目標】 1. Develop English skills for academic and professional purposes

2. Acquire greater vocabulary to communicate in technical English

3. Improve listening skills

4. Enhance the ability to better understand spoken English 【授業計画】

1. Check-up

2. Numbers 1

3. Numbers 2

4. Counting

5. Comparisons 1

6. Comparisons 2

7. Writing 1

8. Writing 2

9. Reports

10. Transition Words 1

11. Transition Words 2

12. Other Scientific/Technical Vocabulary 13. Checking and Confirming

14. Presentations

15. Presentations

【成績評価基準】 The overall grade will be based on class participation, a mid-term,

presentations and homework assignments. 【JABEE 合格】 [JABEE 合格] 【学習教育目標との関連】 [JABEE 関連] 【教科書】”Presenting Science, ” (Second Edition), 2008, Macmillan Language-

House. 【参考書】 [参考資料] 【WEB 頁】 http://homepage3.nifty.com/sumiyo nishiguchi/ 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215990

【対象学生】Nichia STC students and a limited number of students who are

permitted to take this course. 【連絡先】

⇒ カーペンター (共通講義棟 1F 国際連携教育開発センター, 088-656-7643, walter@cicee.tokushima-u.ac.jp) ����

Basic Technical English 1 unit (selection)

Walter Carpenter · Designated Associate Professor / Production Systems Engineering, Department of Mechanical Engineering

Target⟩ This course aims to increase the student’s ability to understand and discuss

professional materials in English. Outline⟩ The general objectives of the course include: 1. Developing students’

reading and listening skills for academic, professional purposes 2. Assisting students’ ability to speak English in real-life situations 3. Helping students develop ability to understand English in various contexts. 4. Increase students’ knowledge of English vocabulary, especially words of a technical and scientific nature.

�������⟩ Technical English, Vocabulary, Scientific Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Technical English”(1.0) Requirement⟩ [要件] Notice⟩ This course is a compulsory one for Nichia STC students. If you are not

a Nichia STC student and want to attend this class, please contact a coordinator in your department for getting permission for your attendance.

Goal⟩

1. Develop English skills for academic and professional purposes

2. Acquire greater vocabulary to communicate in technical English

3. Improve listening skills

4. Enhance the ability to better understand spoken English

Schedule⟩

1. Check-up

2. Numbers 1

3. Numbers 2

4. Counting

5. Comparisons 1

6. Comparisons 2

7. Writing 1

8. Writing 2

9. Reports

10. Transition Words 1

11. Transition Words 2

12. Other Scientific/Technical Vocabulary 13. Checking and Confirming

14. Presentations

15. Presentations

Evaluation Criteria⟩ The overall grade will be based on class participation, a

mid-term, presentations and homework assignments. Jabee Criteria⟩ [JABEE 合格] Relation to Goal⟩ [JABEE 関連] Textbook⟩ ”Presenting Science, ” (Second Edition), 2008, Macmillan Language-

House. Reference⟩ [参考資料] Webpage⟩ http://homepage3.nifty.com/sumiyo nishiguchi/ Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215990

Student⟩ Nichia STC students and a limited number of students who are permitted

to take this course.

Contact⟩

⇒ Carpenter (共通講義棟 1F 国際連携教育開発センター, +81-88-656-7643, walter@cicee.tokushima-u.ac.jp) ����